OpenStudy (anonymous):

What is the radius of a circle with the circumfrece of 6?

OpenStudy (saifoo.khan):

c = pi d

OpenStudy (saifoo.khan):

6/pi =1.9

myininaya (myininaya):

C=2*pi*r 6=2*pi*r 6/(2pi)=r 3/pi=r

OpenStudy (saifoo.khan):

1.9/2 0.95 cm